OpenStudy (anonymous):

find two nonnegative numbers whose product is 50 and whose sum is a minimum?

OpenStudy (amistre64):

integers?

OpenStudy (anonymous):

25 *2

OpenStudy (anonymous):

let the two numbers be x and 50/x

OpenStudy (anonymous):

then form y=x+50/x

OpenStudy (anonymous):

Now differentiate dy/dx=1-50/x^2

OpenStudy (amistre64):

ima gonna guess maybe its sqrt(50) amd sqrt(50) :)

OpenStudy (anonymous):

Equate it with zero which will give you x^2=50

OpenStudy (amistre64):

somewhere in me head I recall that a square has the greatest area and the smallest perimeter values

OpenStudy (anonymous):

so x=sqrt 50 and the another number is sqrt 50 as well

OpenStudy (anonymous):

Yeap amistre

OpenStudy (anonymous):

Who needs calculus...